OpenStudy (anonymous):

how do you solve x over 3 = x over 2 - 2

OpenStudy (anonymous):

\[x/3 = x/2 - 2\] You need to make a common denominator so we'll make it 6 so we get:\[2x/6 =3x/6 - 12/6\] Subtracting the 3x/6 to the other side we get: \[-x/6 = -12/6\] Multiplying both sides by 6 we get -x = -12 so x = 12

OpenStudy (anonymous):

Multiply both sides by 6 leaving you with 2x=3x-12 -x=-12 x=12 (modified)